OpenStudy (anonymous):

How do you factor x^3+8? I know I extract an x and get x(x2+8), but kind of stumped there.

OpenStudy (anonymous):

x(x^2+8) rather

OpenStudy (anonymous):

Convert in the form of a to the power of 3 and 2 to the power of 3 and simplify

OpenStudy (anonymous):

I have no idea what that means.

OpenStudy (anonymous):

someone help me at the end of my problem

OpenStudy (anonymous):

x^3+6 = (x^3+2^3)=(x+2)(x^2-2x+2^2)

OpenStudy (anonymous):

It was x^4+8, not 3

OpenStudy (anonymous):

sorry, x^3+8

OpenStudy (anonymous):

roots are complex